OpenStudy (anonymous):

find the derivative function of y given that y= e^((x^2)(y)) - 100cos(x^2)

OpenStudy (anonymous):

y+100cos(x^2)=e^(x^2y) ln(y+100cos(x^2))=x^2 y 1/y+100cos(x^2) (y'-200xsin(x^2))=2xy+y'x^2 then just simpify